The lucid and lucmath packages^{∗}

Sebastian Rahtz

1997/01/31

1 Packages

1.1 Lucid Package

Set up text fonts for Adobe Lucida.

1 ⟨∗lucid⟩

 2 \renewcommand{\sfdefault}{plcs}
 3 \renewcommand{\rmdefault}{plc}
 4 \renewcommand{\ttdefault}{pcr}

5 ⟨/lucid⟩

1.2 Lucmath Package

Set up math fonts for Adobe Lucida.

6 ⟨∗lucmath⟩

 7 \DeclareSymbolFont{operators}{\encodingdefault}{\rmdefault}{m}{n}
 8 \DeclareSymbolFont{letters}{OML}{plcm}{m}{it}
 9 \DeclareSymbolFont{symbols}{OMS}{plcy}{m}{n}
10 \DeclareSymbolFont{largesymbols}{OMX}{plcv}{m}{n}
11 \SetSymbolFont{letters}{bold}{OML}{plcm}{m}{it}
12 \SetSymbolFont{operators}{bold}{\encodingdefault}{\rmdefault}{b}{n}
13 \SetSymbolFont{operators}{normal}{\encodingdefault}{\rmdefault}{m}{n}

reset the text in math macros

14 \DeclareMathAlphabet      {\mathbf}{\encodingdefault}{\rmdefault}{b}{n}
15 \DeclareMathAlphabet      {\mathrm}{\encodingdefault}{\rmdefault}{m}{n}
16 \DeclareMathAlphabet      {\mathsf}{\encodingdefault}{\sfdefault}{m}{n}
17 \DeclareMathAlphabet      {\mathit}{\encodingdefault}{\rmdefault}{m}{it}
18 \DeclareMathAlphabet      {\mathtt}{\encodingdefault}{\ttdefault}{m}{n}
19 \SetMathAlphabet{\mathbf}{bold}{\encodingdefault}{\rmdefault}{b}{n}
20 \SetMathAlphabet{\mathsf}{bold}{\encodingdefault}{\sfdefault}{b}{n}
21 \SetMathAlphabet{\mathrm}{bold}{\encodingdefault}{\rmdefault}{b}{n}
22 \SetMathAlphabet{\mathit}{bold}{\encodingdefault}{\rmdefault}{b}{it}
23 \SetMathAlphabet{\mathtt}{bold}{\encodingdefault}{\ttdefault}{b}{n}

24 ⟨/lucmath⟩

∗This file has version number v4.0, last revised 1997/01/31.